please answer questions #1 and #3 using descriptive and inferential statistics.

Retail Store1 Retail Store2 Competitor Customers Downtown 600 $35,623 450$33,286 550$29,335 600 $50,295 850$62,891 800$65,203 750$52,465 650$45,206 600$40,789 600$44,924 650 $50,928 1,500$70,678 Customers Customers 480 $26,000 450 $20,325 490 $33,119 540 $39,259 620$44,601 650 $45,112 660 $39,103 590 $35,104 670 $34,119 600 $35,109 700 $39,678 1050$58,471 500$32,679 500$28,345 580$42,892 650 $45,961 700$51,891 730 $52,943 700$44,125 640 $60,200 600 $56,998 650$54.934 700 $50,078 1,200$78,954 Janua Februa ri June August September October November December Total 7,500 $450,000 8,150$600,000 8,600$581,623 1. 2. Using descriptive statistics, what can we learn about customers of each store? Lets make the following assumptions about the customers of each store: Retail #1: p-6096 Retail #2: p-7096 Competitor: p-50% b. c. 3. Using the inferential statistics (assuming these stores can service as samples for the metro city area), assist Mr. Speed in determining (within 95% confidence level) if the customer spending patterns will allow him to meet his $9,000,000 sales by 2023 with 7 stores.
Answer #1

1. Descriptive statistics include measures of central tendency such mean, median, mode and measures of dispersion such as range and std. deviation, etc,.

The following excel image shows the values of various descriptive statistical parameters:

Store 1 customers 480 450 490 540 620 650 660 590 670 600 700 1050 7500 625 610 store 2 customers 500 500 580 650 700 730 700

The store 1 has the least mean number of customers whereas the competitor store has the highest mean number of customers.

700 is the more frequent number(mode) of customers of store 2 and 600 is the more frequent number of customers of competitor store whereas there is no certain frequent number of customers of store1.

The standard deviation of store 1 is least, so the consistency of number customers is more to store 1 whereas for for competitor store the number of customers is less consistent as it has more std. deviation, i.e., spread of customers from the mean number of customers is large.

Now, let us know on an average, how much does a customer of any store pays to that store by puchasing the things:

Store 1: $450,000/7,500 = $60

Store 2: 600,000/8,150 = $74

Store 3: 581,623/8,600 = $68

Thus, store 2 gets more amount per customer and store 1 gets the less amount per customer.